PIECES OF THE PROPELLANE PUZZLE: AN INVESTIGATION OF ROVIBRATIONAL COUPLING

\underline{ROBYNNE KIRKPATRICK}, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, 153 Gilbert Hall, Corvallis, OR 97331-4003; TONY MASIELLO, TOM BLAKE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99532; ALFONS WEBER, National Institute of Standards and Technology, Gaithersburg, MD 20899; ART MAKI, Retired, National Institute of Standards and Technology.

[1.1.1]Propellane (C_{5}H_{6}) is a small ring D_{3h} molecule with an axial bond that continues to be a topic of debate. Recently propellane was subjected to high resolution spectroscopic scrutiny for the first time, and its ν,l,J,K structure is being determined. It was found that ν_{15}, the axial antisymmetric C-C stretching mode that has a remarkably high dipole derivative, and the ν_{12} H wagging modes are rotationally coupled.